OpenStudy (anonymous):

suppose that at the end of any given year an automobile is worth only 80% of its value at the beginning of that year. If a car cost $29000 new, what is its value at the end of five years

OpenStudy (anonymous):

Answer is $ 9502.72 do u want the working??

OpenStudy (anonymous):

i tried that answer is was not correct

OpenStudy (anonymous):

\[29000\times0.8^{5}=$9502.72\]

OpenStudy (anonymous):

here we can use the compounding formula as follows : ( R )^n ( 100 - R )^n A = P (1 - ----) = P (--------) ( 100) ( 100 ) here P = 29000, R = 20, n =5 so ( 100 - R )^n (100 - 20 )^5 ( 80 )^5 8^5 A = P (--------) = 29000 (--------) = 29000 (----- ) = 29000 * ----- ( 100) ( 100 ) ( 100 ) 10^5 32768 = 29000 * -------- = 9502.72 100000 So the car's value after 5 years will be $ 9502.72

OpenStudy (anonymous):

I have given the full working.......... How did you check the answer??? From the answer key??? Sometimes there can be printing errors......